### v2

# Robust MRI Reconstruction by Smoothed Unrolling (SMUG)

Abstract
As the popularity of deep learning (DL) in the field of magnetic resonance imaging (MRI) continues to rise, recent research has indicated that DL-based MRI reconstruction models might be excessively sensitive to minor input disturbances, including worst-case or random additive perturbations. This sensitivity often leads to unstable aliased images. This raises the question of how to devise DL techniques for MRI reconstruction that can be robust to these variations. To address this problem, we propose a novel image reconstruction framework, termed Smoothed Unrolling (SMUG), which advances a deep unrolling-based MRI reconstruction model using a randomized smoothing (RS)-based robust learning approach. RS, which improves the tolerance of a model against input noise, has been widely used in the design of adversarial defense approaches for image classification tasks. Yet, we find that the conventional design that applies RS to the entire DL-based MRI model is ineffective. In this paper, we show that SMUG and its variants address the above issue by customizing the RS process based on the unrolling architecture of DL-based MRI reconstruction models. We theoretically analyze the robustness of our method in the presence of perturbations. Compared to vanilla RS and other recent approaches, we show that SMUG improves the robustness of MRI reconstruction with respect to a diverse set of instability sources, including worst-case and random noise perturbations to input measurements, varying measurement sampling rates, and different numbers of unrolling steps. Our code is available at https://github.com/sjames40/SMUG_journal.
Index Terms:
Magnetic resonance imaging, machine learning, deep unrolling, robustness, randomized smoothing, compressed sensing.
I Introduction
Magnetic resonance imaging (MRI) is a popular noninvasive imaging modality, which involves a sequential and slow data collection. As such, MRI scans can be accelerated by collecting limited data. In this case, the process of image reconstruction requires tackling an ill-posed inverse problem. To deliver accurate image reconstructions from such limited information, compressed sensing (CS) [1] has been extensively used. Conventional CS-MRI assumes the underlying image’s sparsity (in practice, enforced in the wavelet domain [2] or via total variation [3]). As further improvement to conventional CS, various learned sparse signal models have been well-studied, such as involving patch-based synthesis dictionaries [4, 5], or sparsifying transforms [6, 7]. Learned transforms have been shown to offer an efficient and effective framework for sparse modeling in MRI [8].
Recently, due to the outstanding representation power of convolutional neural networks (CNNs), they have been applied in single-modality medical imaging synthesis and reconstruction  [9, 10, 11, 12]. The U-Net architecture, presented in [13] and used in several studies, is a popular deep CNN for many tasks involving image processing. They exhibit two key features: the use of a diminishing path for gathering contextual information, and a symmetric expansion path for precise localization.
Hybrid-domain DL-based image reconstruction methods, such as Model-based reconstruction using Deep Learned priors (MoDL[11], Iterative Shrinkage-Thresholding Algorithm (ISTA-Net) [14], etc., are used to enhance stability and performance by ensuring data consistency in the training and reconstruction phases. In MR imaging, data consistency layers are often essential in reconstruction networks to ensure the image agrees with the measurement model [15, 16]. Various methods such as [17, 18, 14, 11] maintain this consistency by deep unrolling-based architectures, which mimic a traditional iterative algorithm and learn the associated regularization parameters. Other approaches ensure data consistency by applying methods such as denoising regularization [19] and plug-and-play techniques [20]. Despite their recent advancements, DL-based MRI reconstruction models are shown to be vulnerable to tiny changes or noise in the input, shifts in the measurement sampling rate [21, 22], and varying iteration numbers in unrolling schemes [23]. In such cases, the resulting images from DL models are of inferior quality which could possibly lead to inaccurate diagnoses and, consequently, undesirable clinical consequences.
It is of much importance in medical imaging applications to learn reconstruction models that are robust to various measurement artifacts, noise, and scan or data variations at test time. Although there exist numerous robustification techniques  [24, 25, 26, 27] to tackle the instability of DL models in image classification tasks, methods to enhance the robustness of DL-based MRI reconstruction models are less explored due to their regression-based learning targets. Methods such as randomized smoothing (RS) and its variations [26, 27, 28], are often used in image classification. They diverge from traditional defense methods [24, 25] such as adversarial training, which provide some empirical robustness but are computationally expensive and could fail under more diverse perturbations. RS ensures the model’s stability within a radius surrounding the input image [26], which could be critical for medical use cases such as MRI. Recent early-stage research has begun to apply RS to DL-based MRI reconstruction in an end-to-end manner [29]. However, the end-to-end RS approach might not always be an appropriate fit for all image reconstructors, such as physics-based and hybrid methods.
In our recent conference work [30], we proposed integrating the RS approach within the MoDL framework for the problem of MR image reconstruction. This is accomplished by using RS in each unrolling step and at the intermediate unrolled denoisers in MoDL. This strategy is underpinned by the ‘pre-training + fine-tuning’ technique [31, 27]. In [30], we empirically showed that this approach is effective. In this paper, we provide an analysis and conditions under which the proposed smoothed unrolling (SMUG) technique is robust against perturbations. Furthermore, we introduce a novel weighted smoothed unrolling scheme that learns image-wise weights during smoothing unlike conventional RS. This approach further improves the reconstruction performance. Furthermore, in this work, we evaluate worst-case additive perturbations in k-space or measurement space, in contrast to [30], where image-space perturbations were considered. Moreover, this paper includes additional comprehensive experimental results and application of our framework to multiple deep reconstruction models.
I-A Contributions
The main contributions of this work are as follows:
  • We propose SMUG that systematically integrates RS in each iteration within deep unrolled reconstruction.
  • We provide a theoretical analysis to demonstrate the robustness of SMUG.
  • We enhance the performance of SMUG by introducing weighted smoothing as an improvement over conventional RS and showcase the resulting gains.
  • We integrate the technique into multiple unrolled models including MoDL [11] and ISTA-Net [14] and demonstrate improved robustness of our methods compared to the original schemes. We also show advantages for SMUG over end-to-end RS [29], Adversarial Training (AT) [32], Deep Equilibrium (Deep-Eq) models [33], and a leading diffusion-based model [34]. Extensive experiments demonstrate the potential of our approach in handling various types of reconstruction instabilities.
I-B Paper Organization
The remainder of the paper is organized as follows. In Section II, we present preliminaries and the problem statement. Our proposed method is described in Section III. Section IV presents experimental results and comparisons, and we conclude in Section V.
II Preliminaries and Problem Statement
II-A Setup of MRI Reconstruction
Many medical imaging approaches involve ill-posed inverse problems such as the work in [35], where the aim is to reconstruct the original signal $\mathbf{x}\in\mathbb{C}^{q}$ (vectorized image) from undersampled k-space (Fourier domain) measurements $\mathbf{y}\in\mathbb{C}^{p}$ with $p<q$. The imaging system in MRI can be modeled as a linear system $\mathbf{y}\approx\mathbf{A}\mathbf{x}$, where $\mathbf{A}$ may take on different forms for single-coil or parallel (multi-coil) MRI, etc. For example, in the single coil Cartesian MRI acquisition setting, $\mathbf{A}=\mathbf{M}\mathbf{F}$, where $\mathbf{F}$ is the 2D discrete Fourier transform and $\mathbf{M}$ is a masking operator that implements undersampling. With the linear observation model, MRI reconstruction is often formulated as
$\hat{\mathbf{x}}=\underset{\mathbf{x}}{\arg\min}~{}\|\mathbf{A}\mathbf{x}-\mathbf{y}\|^{2}_{2}+\lambda\mathcal{R}(\mathbf{x}),$ (1)
where $\mathcal{R}(\cdot)$ is a regularization function (e.g., $\ell_{1}$ norm in the wavelet domain to impose a sparsity prior [2]), and $\lambda>0$ is the regularization parameter.
MoDL [36] is a recent popular supervised deep learning approach inspired by the MR image reconstruction optimization problem in (1). MoDL combines a denoising network with a data-consistency (DC) module in each iteration of an unrolled architecture. In MoDL, the hand-crafted regularizer, $\mathcal{R}$, is replaced by a learned network-based prior $\left\|\mathbf{x}-\mathcal{D}_{\bm{\theta}}(\mathbf{x})\right\|_{2}^{2}$ involving a network $\mathcal{D}_{\bm{\theta}}$. MoDL attempts to optimize this loss by initializing $\mathbf{x}^{0}=\mathbf{A}^{H}\mathbf{y}$, and then iterating the following process for a number of unrolling steps indexed by $n\in\{0,\dots,N-1\}$. Specifically, MoDL iterations are given by
$\mathbf{x}^{n+1}=\operatorname*{arg\,min}_{\mathbf{x}}\|\mathbf{A}\mathbf{x}-\mathbf{y}\|_{2}^{2}+\lambda\|\mathbf{x}-\mathcal{D}_{\bm{\theta}}(\mathbf{x}^{n})\|_{2}^{2}.$ (2)
After $N$ iterations, we denote the final output of MoDL as $\mathbf{x}^{N}=\bm{F}_{\text{MoDL}}(\mathbf{x}^{0})$. The weights of the denoiser are shared across the $N$ blocks and are learned in an end-to-end supervised manner [11].
II-B Lack of Robustness of DL-based Reconstructors
In [21], it was demonstrated that deep learning-based MRI reconstruction can exhibit instability, when confronted with subtle, nearly imperceptible input perturbations. These perturbations are commonly referred to as ‘adversarial perturbations’ and have been extensively investigated in the context of DL-based image classification tasks, as outlined in [37]. In the context of MRI, these perturbations represent the worst-case additive perturbations, which can be used to evaluate method sensitivity and robustness [21, 38]. Let $\bm{\delta}$ denote a small perturbation of the measurements that falls in an $\ell_{\infty}$ ball of radius $\epsilon$, i.e., $\|\bm{\delta}\|_{\infty}\leq\epsilon$. Adversarial disturbances then correspond to the worst-case input perturbation vector $\bm{\delta}$ that maximizes the reconstruction error, i.e.,
$\begin{array}[]{ll}\max_{\|\bm{\delta}\|_{\infty}\leq\epsilon}\|\bm{F}_{\text{{{MoDL}}}}(\mathbf{A}^{H}(\mathbf{y}+\bm{\delta}))-\mathbf{t}\|_{2}^{2},\end{array}$ (3)
where $\mathbf{t}$ is a ground truth target image from the training set (i.e., label). The operator $\mathbf{A}^{H}$ transforms the measurements $\mathbf{y}$ to the image domain, and $\mathbf{A}^{H}\mathbf{y}$ is the input (aliased) image to the reconstruction model. The optimization problem in (3) can be effectively solved using the iterative projected gradient descent (PGD) method [24].
In Fig. 1-(a) and (b), we show reconstructed images using MoDL originating from a benign (i.e., undisturbed) input and a PGD scheme-perturbed input, respectively. It is evident that the worst-case input disturbance significantly deteriorates the quality of the reconstructed image. While one focus of this work is to enhance robustness against input perturbations, Fig.1-(c) and (d) highlight two additional potential sources of instability that the reconstructor (MoDL) can encounter during testing: variations in the measurement sampling rate (resulting in “perturbations” to the sparsity of the sampling mask in $\mathbf{A}$) [21], and changes in the number of unrolling steps [23]. In scenarios where the sampling mask (Fig.1-(c)) or number of unrolling steps (Fig.1-(d)) deviate from the settings used during MoDL training, we observe a significant degradation in performance compared to the original setup (Fig.1-(a)), even in the absence of additive measurement perturbations. In Section IV, we demonstrate how our method improves the reconstruction robustness in the presence of different types of perturbations, including those in Fig.1.
Figure 1: MoDL’s instabilities resulting from perturbations to input data, the measurement sampling rate, and the number of unrolling steps used at testing phase shown on an image from the fastMRI dataset [39]. We refer readers to Section IV for further details about the experimental settings. (a) MoDL reconstruction from benign (i.e., without additional noise/perturbation) measurements with 4×4$\times$4 × acceleration (i.e., 25% sampling rate) and 8 unrolling steps. (b) MoDL reconstruction from disturbed input with perturbation strength $\epsilon=0.02$ (see Section IV-A). (c) MoDL reconstruction from clean measurements with 2×2$\times$2 × acceleration (i.e., 50% sampling), and using 8 unrolling steps. (d) MoDL reconstruction from clean or unperturbed measurements with 4×4$\times$4 × acceleration and 16 unrolling steps. In (b), (c), and (d), the network trained in (a) is used.
II-C Randomized Smoothing (RS)
Randomized smoothing, introduced in [26], enhances the robustness of DL models against noisy inputs. It is implemented by generating multiple randomly modified versions of the input data and subsequently calculating an averaged output from this diverse set of inputs.
Given some function $f(\mathbf{x})$, RS formally replaces $f$ with a smoothed version
$g(\mathbf{x})\mathrel{\mathop{:}}=\mathbb{E}_{\bm{\eta}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})}[f(\mathbf{x}+\bm{\eta})]\>,$ (4)
where $\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ denotes a Gaussian distribution with zero mean and element-wise variance $\sigma^{2}$, and $\mathbf{I}$ denotes the identity matrix of appropriate size. Prior research has shown that RS has been effective as an adversarial defense approach in DL-based image classification tasks [26, 27, 40]. However, the question of whether RS can significantly improve the robustness of MoDL and other image reconstructors has not been thoroughly explored. A preliminary investigation in this area was conducted by [29], which demonstrated the integration of RS into MR image reconstruction in an end-to-end (E2E) setting. We can formulate image reconstruction using RS-E2E as
$\mathbf{x}_{\textsc{RS-E2E}}=\mathbb{E}_{\bm{\eta}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})}[\bm{F}_{\text{{MoDL}}}(\mathbf{A}^{H}(\mathbf{y}+\bm{\eta}))].$ (RS-E2E)
This formulation aligns with the one used in [29], where the random noise vector $\bm{\eta}$ is directly added to $\mathbf{y}$ in the frequency domain (complex-valued), followed by multiplication with $\mathbf{A}^{H}$ to obtain the input image for MoDL. The noisy measurements are also utilized in each iteration in MoDL. RS-E2E can be identically formulated for alternative reconstruction models.
Fig. 2 shows a block diagram of RS-E2E-backed MoDL. This RS-integrated MoDL is trained with supervision in the standard manner. Although RS-E2E represents a straightforward application of RS to MoDL, it remains unclear if the formulation in (RS-E2E) is the most effective method to incorporate RS into unrolled algorithms such as MoDL, considering the latter’s specialties, e.g., the involved denoising and the data-consistency (DC) steps.
As such, for the rest of the paper, we focus on studying the following questions (Q1)(Q4).
(Q1): How should RS be integrated into an unrolled algorithm such as MoDL? (Q2): How do we learn the network $\mathcal{D}_{\bm{\theta}}(\cdot)$ in the presence of RS operations? (Q3): Can we prove the robustness of SMUG in the presence of data perturbations? (Q4): Can we further improve the RS operation in SMUG for enhanced image quality or sharpness?
III Methodology
In this section, we address questions (Q1)(Q4) by taking the unrolling characteristics of MoDL into the design of an RS-based MRI reconstruction. The proposed novel integration of RS with MoDL is termed Smoothed Unrolling (SMUG). We also explore an extension to SMUG. We note that while we develop our methods based on MoDL, in the last subsection, we discuss incorporating our approaches within other unrolling methods such as ISTA-Net.
III-A Solution to (Q1): RS at intermediate unrolled denoisers
As illustrated in Fig.,2, the RS operation in RS-E2E is typically applied to MoDL in an end-to-end manner. This does not shed light on which component of MoDL needs to be made more robust. Here, we explore integrating RS at each intermediate unrolling step of MoDL. In this subsection, we present SMUG, which applies RS to the denoising network. This seemingly simple modification is related to a robustness certification technique known as “denoised smoothing” [27]. In this technique, a smoothed denoiser is used, proving to be sufficient for establishing robustness in the model. We use $\mathbf{x}_{\textrm{S}}^{n}$ to denote the $n$-th iteration of SMUG. Starting from $\mathbf{x}_{\textrm{S}}^{0}=\mathbf{A}^{H}\mathbf{y}$, the procedure is given by
$\mathbf{x}_{\textrm{S}}^{n+1}=\operatorname*{arg\,min}_{\mathbf{x}}\|\mathbf{A}\mathbf{x}-\mathbf{y}\|_{2}^{2}+\lambda\|\mathbf{x}-\mathbb{E}_{\bm{\eta}}\big{[}\mathcal{D}_{\bm{\theta}}(\mathbf{x}_{\textrm{S}}^{n}+\bm{\eta})\big{]}\|_{2}^{2}\>,$ (5)
where $\bm{\eta}$ is drawn from $\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$. After $N-1$ iterations, the final output of SMUG is denoted by $\mathbf{x}^{N}_{\text{S}}=\bm{F}_{\text{SMUG}}(\mathbf{x}^{0})$. Fig. 3 presents the architecture of SMUG.
Figure 2: A schematic overview of RS-E2E. Here, iterative unrolling takes place between the data consistency and denoising blocks for multiple noisy versions of the input.
Figure 3: Architecture of SMUG. Here, for every unrolling step, after applying the denoiser on each noisy version of the input, the data consistency is applied on the average of the denoised images.
Figure 4: Architecture of weighted SMUG. Here, we extend SMUG by including the weight encoder and the use of weighted randomized smoothing.
III-B Solution to (Q2): SMUG’s pre-training & fine-tuning
In this subsection, we develop the training scheme of SMUG. Inspired by the currently celebrated “pre-training + fine-tuning” technique [31, 27], we propose to train SMUG following this learning paradigm. Our rationale is that pre-training can provide a robustness-aware initialization of the DL-based denoising network for fine-tuning. To pre-train the denoising network $\mathcal{D}_{\bm{\theta}}$, we consider a mean squared error (MSE) loss that measures the Euclidean distance between images denoised by $\mathcal{D}_{\bm{\theta}}$ and the target (ground truth) images, denoted by $\mathbf{t}$. This leads to the pre-training step:
$\bm{\theta}_{\mathrm{pre}}=\displaystyle\operatorname*{arg\,min}_{\bm{\theta}}\mathbb{E}_{\mathbf{t}\in\mathcal{T}}[\mathbb{E}_{\bm{\eta}}||\mathcal{D}_{\bm{\theta}}(\mathbf{t}+\bm{\eta})-\mathbf{t}||_{2}^{2}]\>,$ (6)
where $\mathcal{T}$ is the set of ground truth images in the training dataset. Next, we develop the fine-tuning scheme to improve $\bm{\theta}_{\mathrm{pre}}$ based on the labeled/paired MRI dataset. Since RS in SMUG (Fig. 3) is applied to every unrolling step, we propose an unrolled stability (UStab) loss for fine-tuning $\mathcal{D}_{\bm{\theta}}$:
$\ell_{\mathrm{UStab}}(\bm{\theta};\mathbf{y},\mathbf{t})=\sum_{n=0}^{N-1}\mathbb{E}_{\bm{\eta}}||\mathcal{D}_{\bm{\theta}}(\mathbf{x}^{n}+\bm{\eta})-\mathcal{D}_{\bm{\theta}}(\mathbf{t})||^{2}_{2}\>.$ (7)
The UStab loss in (7) relies on the target images, bringing in a key benefit: the denoising stability is guided by the reconstruction accuracy of the ground-truth image, yielding a graceful trade-off between robustness and accuracy.
Integrating the UStab loss, defined in (7), with the standard reconstruction loss, we obtain the fine-tuned $\bm{\theta}$ by minimizing $\mathbb{E}_{(\mathbf{y},\mathbf{t})}[\ell(\bm{\theta};\mathbf{y},\mathbf{t})]$, where
$\ell(\bm{\theta};\mathbf{y},\mathbf{t})=\ell_{\mathrm{UStab}}(\bm{\theta};\mathbf{y},\mathbf{t})+\lambda_{\ell}\|\bm{F}_{\text{SMUG}}(\mathbf{A}^{H}\mathbf{y})-\mathbf{t}\|_{2}^{2},$ (8)
with $\lambda_{\ell}>0$ representing a regularization parameter to strike a balance between the reconstruction error (for accuracy) and the denoising stability (for robustness) terms. We initialize $\bm{\theta}$ as $\bm{\theta}_{\mathrm{pre}}$ when optimizing (8) using standard optimizers such as Adam [41].
In practice, the same dataset is used for fine-tuning as pre-training because the pre-trained model is initially trained solely as a denoiser, while the fine-tuning process aims at integrating the entire regularization strategy applied to the MoDL framework. This approach ensures that the fine-tuning optimally adapts the model to the specific enhancements introduced by our robustification strategies.
III-C Answer to (Q3): Analyzing the robustness of SMUG in the presence of data perturbations
The following theorem discusses the robustness (i.e., sensitivity to input perturbations) achieved with SMUG. Note that all norms on vectors (resp. matrices) denote the $\ell_{2}$ norm (resp. spectral norm) unless indicated otherwise.
Theorem 1.
Assume the denoiser network’s output is bounded in norm. Given the initial input image $\mathbf{A}^{H}\mathbf{y}$ obtained from measurements $\mathbf{y}$, let the SMUG reconstructed image at the $n$-th unrolling step be $\mathbf{x}^{n}_{\text{S}}(\mathbf{A}^{H}\mathbf{y})$ with RS variance of $\sigma^{2}$. Let $\bm{\delta}$ denote an additive perturbation to the measurements $\mathbf{y}$. Then,
$\|\mathbf{x}^{n}_{\text{S}}(\mathbf{A}^{H}\mathbf{y})-\mathbf{x}^{n}_{\text{S}}(\mathbf{A}^{H}(\mathbf{y}+\bm{\delta}))\|\leq C_{n}\|\bm{\delta}\|,$ (9)
where $C_{n}=\alpha\|\mathbf{A}\|_{2}\begin{pmatrix}\frac{1-\begin{pmatrix}\frac{M\alpha}{\sqrt{2\pi}\sigma}\end{pmatrix}^{n}}{1-\frac{M\alpha}{\sqrt{2\pi}\sigma}}\end{pmatrix}+\|\mathbf{A}\|_{2}\begin{pmatrix}\frac{M\alpha}{\sqrt{2\pi}\sigma}\end{pmatrix}^{n}$, with $\alpha=\|(\mathbf{A}^{H}\mathbf{A}+\mathbf{I})^{-1}\|_{2}$ and $M=2\max_{\mathbf{x}}(\|\mathcal{D}_{\bm{\theta}}(\mathbf{x})\|)$.
The proof is provided in the Appendix. Note that the output of SMUG $\mathbf{x}^{n}_{\text{S}}(\cdot)$ depends on both the initial input (here $\mathbf{A}^{H}\mathbf{y}$) and the measurements $\mathbf{y}$. We abbreviated it to $\mathbf{x}^{n}_{\text{S}}(\mathbf{A}^{H}\mathbf{y})$ in the theorem and proof for notational simplicity. The constant $C_{n}$ depends on the number of iterations or unrolling steps $n$ as well as the RS standard deviation parameter $\sigma$. For large $\sigma$, the robustness error bound for SMUG clearly decreases as the number of iterations $n$ increases. In particular, if $\sigma>M\alpha/\sqrt{2\pi}$, then as $n\rightarrow\infty$, $C_{n}\rightarrow\alpha\left\|\mathbf{A}\right\|_{2}/\begin{pmatrix}1-\frac{M\alpha}{\sqrt{2\pi\sigma}}\end{pmatrix}$. Furthermore, as $\sigma\to\infty$, $C_{n}\to C\triangleq\alpha\left\|\mathbf{A}\right\|_{2}$. Clearly, if $\alpha\leq 1$ and $\left\|\mathbf{A}\right\|_{2}\leq 1$ (normalized), then $C\leq 1$.
Thus, for sufficient smoothing, the error introduced in the SMUG output due to input perturbation never gets worse than the size of the input perturbation. Therefore, the output is stable with respect to (w.r.t.) perturbations. These results corroborate experimental results in Section IV on how SMUG is robust (whereas other methods, such as vanilla MoDL, breakdown) when increasing the number of unrolling steps at test time, and is also more robust for larger $\sigma$ (with good accuracy-robustness trade-off).
The only assumption in our analysis is that the denoiser network output is bounded in norm. This consideration is handled readily when the denoiser network incorporates bounded activation functions such as the sigmoid or hyperbolic tangent. Alternatively, if we expect image intensities to lie within a certain range, a simple clipping operation in the network output would ensure boundedness for the analysis.
A key distinction between SMUG and prior works, such as RS-E2E [29], is that smoothing is performed in every iteration. Moreover, while [29] assumes the end-to-end mapping is bounded, in MoDL or SMUG, it clearly isn’t because the data-consistency step’s output is unbounded as $\mathbf{y}$ grows.
We remark that our intention with Theorem 1 is to establish a baseline of robustness intrinsic to models with unrolling architectures.
III-D Solution to (Q4): Weighted Smoothing
In this subsection, we present a modified formulation of randomized smoothing to improve its performance in SMUG. Randomized smoothing in practice involves uniformly averaging images denoised with random perturbations. This can be viewed as a type of mean filter, which leads to oversmoothing of structural information in practice. As such, we propose weighted randomized smoothing, which employs an encoder to assess a weighting (scalar) for each denoised image and subsequently applies the optimal weightings while aggregating images to enhance the reconstruction performance. Our method not only surpasses the SMUG technique but also excels in enhancing image sharpness across various types of perturbation sources. This allows for a more versatile or flexible and effective approach for improving image quality under different conditions.
The weighted randomized smoothing operation applied on a function $f(\cdot)$ is as follows:
$g_{\textrm{w}}(\mathbf{x}):=\frac{\mathbb{E}_{\bm{\eta}}[w(\mathbf{x}+\bm{\eta})f(\mathbf{x}+\bm{\eta})]}{\mathbb{E}_{\bm{\eta}}[w(\mathbf{x}+\bm{\eta})]}\>,$ (10)
where $w(\cdot)$ is an input-dependent weighting function.
Based on the weighted smoothing in (10), we introduce Weighted SMUG. This approach involves applying weighted RS at each denoising step, and the weighting encoder is trained in conjunction with the denoiser during the fine-tuning stage. For the weighting encoder in our experiments, we use a simple architecture consisting of five successive convolution, batch normalization, and ReLU activation layers followed by a linear layer and Sigmoid activation. Specifically, in the $n$-th unrolling step, we use a weighting encoder $\mathcal{E}_{\bm{\phi}}$, parameterized by $\bm{\phi}$, to learn the weight of each image used for (weighted) averaging. Here, we use $\mathbf{x}^{n}_{\textrm{W}}$ to denote the output of the $n$-th block. Initializing $\mathbf{x}^{0}_{\textrm{W}}=\mathbf{A}^{H}\mathbf{y}$, the output of Weighted SMUG w.r.t. $n$ is
$\begin{split}&\mathbf{x}_{\textrm{W}}^{n+1}=\operatorname*{arg\,min}_{\mathbf{x}}~{}\|\mathbf{A}\mathbf{x}-\mathbf{y}\|_{2}^{2}~{}+\\ &\lambda\begin{Vmatrix}\mathbf{x}-\frac{\mathbb{E}_{\bm{\eta}}[\mathcal{E}_{\bm{\phi}}(\mathbf{x}_{\textrm{W}}^{n}+\bm{\eta})\mathcal{D}_{\bm{\theta}}(\mathbf{x}_{\textrm{W}}^{n}+\bm{\eta})]}{\mathbb{E}_{\bm{\eta}}[\mathcal{E}_{\bm{\phi}}(\mathbf{x}_{\textrm{W}}^{n}+\bm{\eta})]}\end{Vmatrix}_{2}^{2}\>.\end{split}$ (11)
After $N$ iterations, the final output of Weighted SMUG is $\mathbf{x}^{N}_{\text{W}}=\bm{F}_{\text{wSMUG}}(\mathbf{x}^{0})$. Figure 4 illustrates the block diagram of weighted SMUG.
Furthermore, we extend the “pre-training + fine-tuning” approach proposed in Section III-B to the Weighted SMUG method. In this case, we obtain the fine-tuned $\bm{\theta}$ and $\bm{\phi}$ by using
$\min_{\bm{\theta},\bm{\phi}}\,\mathbb{E}_{(\mathbf{y},\mathbf{t})}[\lambda_{l}\|{F}_{\text{wSMUG}}(\mathbf{A}^{H}\mathbf{y})-\mathbf{t}\|_{2}^{2}+\ell_{\mathrm{UStab}}(\bm{\theta};\mathbf{y},\mathbf{t})].$ (12)
III-E Integrating RS into Other Unrolled Networks
In this subsection, we further discuss the extension of our SMUG schemes for other unrolling based reconstructors, using ISTA-Net [14] as an example. The goal is to demonstrate the generality of our proposed approaches for deep unrolled models.
ISTA-Net uses a training loss function composed of discrepancy and constraint terms. In particular, it performs the following for $N$ unrolling steps:
$\bm{r}^{n}=\mathbf{x}^{n-1}-\lambda^{(n)}\mathbf{A}^{H}(\mathbf{A}\mathbf{x}^{n-1}-\mathbf{y})$ (13)
[…]
Similar to SMUG for MoDL, we integrate RS into the network-based regularization (denoising) component of ISTA-Net. This results in the following modification to (14):
[…]
where $\bm{\eta}$ is drawn from $\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$. For weighted SMUG, (14) becomes
[…]
IV Experiments
IV-A Experimental Setup
IV-A1 Models & Sampling Masks
For the MoDL architecture, we use the recent state-of-the-art denoising network Deep iterative Down Network, which consists of $3$ down-up blocks (DUBs) and 64 channels [42]. Additionally, for MoDL, we use $N=8$ unrolling steps with denoising regularization parameter $\lambda=1$. The conjugate gradient method [36], with a tolerance level of $10^{-6}$, is utilized to execute the DC block. We used variable density Cartesian random undersampling masks in k-space, one for each undersampling factor that include a fully-sampled central k-space region and the remaining phase encode lines were sampled uniformly at random. The coil sensitivity maps for all scenarios were generated with the BART toolbox [43]. Extension to the ISTA-Net model is discussed in Section IV-G.
IV-A2 Baselines
We consider two robustification approaches: first is the RS-E2E method [38] presented in (RS-E2E), and the second is Adversarial Training (AT) [32]. Furthermore, we consider other recent reconstruction models, specifically, the Deep Equilibrium (Deep-Eq) method [33] and a leading diffusion-based MRI reconstruction model from [34], which we denote as Score-MRI.
IV-A3 Datasets & Training
For our study, we execute two experimental cases. For the first case, we utilize the fastMRI knee dataset, with 32 scans for validation and 64 unseen scans/slices for testing. In the second case, we employ our method for the fastMRI brain dataset. We used $3000$ training scans in both cases. The k-space data are normalized so that the real and imaginary components are in the range $[-1,1]$. We use a batch size of $2$ and $60$ training epochs. The experiments are run using two A5000 GPUs. The ADAM optimizer [44] is utilized for training the network weights with momentum parameters of $(0.5,0.999)$ and learning rate of $10^{-4}$. The stability parameter $\lambda_{\ell}$ in (8) (and (12)) is tuned so that the standard accuracy of the learned model is comparable to vanilla MoDL. For RS-E2E, we set the standard deviation of Gaussian noise to $\sigma=0.01$, and use $10$ Monte Carlo samplings to implement the smoothing operation. Note that in our experiments, Gaussian noise and corruptions are added to real and imaginary parts of the data with the indicated $\sigma$. For AT, we implemented a 30-step PGD procedure within its minimax formulation with $\epsilon=0.02$. For Score MRI, we used 150 steps for the reverse diffusion process with the pre-trained model in https://github.com/HJ-harry/score-MRI. We fine-tuned a pre-trained Deep-Eq model [^1] https://github.com/dgilton/deep_equilibrium_inverse with the same data as the proposed schemes. Unless specified, training parameters were similar across the compared methods.
IV-A4 Testing
We evaluate our methods on clean data (without additional perturbations), data with randomly injected noise, and data contaminated with worst-case additive perturbations. The worst-case disturbances allow us to see worst-case method sensitivity and are generated by the $\ell_{\infty}$-norm based PGD scheme with 10 steps [21] corresponding to $\left\|\bm{\delta}\right\|_{\infty}\leq\epsilon$, where $\epsilon$ is set nominally as the maximum underlying k-space real and imaginary part magnitude scaled by $0.05$. We will indicate the scaling for $\epsilon$ (e.g., $0.05$) in the results and plots that follow. The quality of reconstructed images is measured using peak signal-to-noise ratio (PSNR) and structure similarity index measure (SSIM)[45]. In addition to the worst-case perturbations and random noise, we evaluate the performance of our methods in the presence of additional instability sources such as (i) different undersampling rates, and (ii) different numbers of unrolling steps.
IV-B Robustness with Additive Perturbations
Figure 5: Reconstruction accuracy box plots for the fastMRI brain dataset with 4x acceleration factor. The additive random Gaussian noise of the second column plots is obtained using standard deviation of $0.01$. The worst-case additive noise of the third column is obtained using the PGD method with $\epsilon=0.02$.
Ground Truth Vanilla MoDL RS-E2E SMUG
Refer to caption Refer to caption Refer to caption Refer to caption
AT Score-MRI Deep Equilibrium Weighted-SMUG
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Visualization of ground truth and reconstructed images using different methods for 4x k-space undersampling, evaluated on PGD-generated worst-case inputs of perturbation strength $\epsilon=0.02$.
Figure 7: Reconstruction accuracy box plots for the fastMRI knee dataset with 4x Acceleration factor. The additive random Gaussian noise of the second column plots is obtained using a standard deviation of $0.01$. The worst-case additive noise of the third column is obtained using the PGD method with $\epsilon=0.02$.
Figure 8: Reconstruction accuracy box plots for the fastMRI knee dataset with 8x Acceleration factor. The additive random Gaussian noise in the second column plots is obtained using a standard deviation of $0.01$. The worst-case additive noise in the third column is obtained using the PGD method with $\epsilon=0.02$.
Ground Truth Vanilla MoDL RS-E2E SMUG
Refer to caption Refer to caption Refer to caption Refer to caption
AT Score-MRI Deep Equilibrium Weighted-SMUG
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Visualization of ground-truth and reconstructed images using different methods for 4x k-space undersampling, evaluated on PGD-generated worst-case inputs of perturbation strength $\epsilon=0.02$.
Ground Truth Vanilla MoDL RS-E2E SMUG
Refer to caption Refer to caption Refer to caption Refer to caption
AT Score-MRI Deep Equilibrium Weighted-SMUG
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Visualization of ground truth and reconstructed images using different methods for 8x k-space undersampling, evaluated on PGD-generated worst-case inputs of perturbation scaling $\epsilon=0.02$.
In this subsection, we present the robustness results of the proposed approaches w.r.t. additive noise. In particular, the evaluation is conducted on the clean, noisy (with added Gaussian noise), and worst-case perturbed (using PGD for each method) measurements. Fig. 5 presents testing set PSNR and SSIM values as box plots for different smoothing architectures, along with vanilla MoDL and the other baselines using the brain dataset. The clean accuracies of Weighted SMUG and SMUG are similar to vanilla MoDL indicating a good clean accuracy vs. robustness trade-off. As indicated by the PSNR and SSIM values, we observe that weighted SMUG, on average, outperforms all other baselines in robust accuracy (the second and third set of box plots of the two rows in Fig. 5). This observation is consistent with the visualization of reconstructed images for the brain dataset in Fig. 6. We note that weighted SMUG requires longer time for training, which represents a trade-off. When comparing to AT, we observe that AT is comparable to SMUG in the case of robust (or worst-case noise) accuracy. However, the drop in clean accuracy (without perturbations) for AT is significantly larger than for SMUG. Furthermore, AT takes a much longer training time as it requires to solve an optimization problem (PGD) for every training data sample at every iteration to obtain the worst-case perturbations. Furthermore, we observe that its effectiveness is degraded for other perturbations including random noise as well as modified sampling rates shown in the next subsection. Importantly, the proposed SMUG and Weighted SMUG are not trained to be robust to any specific perturbations or instabilities, but are nevertheless effective for several scenarios.
In comparison to the diffusion based Score-MRI, the proposed methods perform better in terms of both clean accuracy and random noise accuracy. Although for worst-case perturbations, the PSNR values of Score-MRI are only slightly worse than SMUG, it is important to note that not only the training of diffusion-based models takes longer than our method, but also the inference time is longer as Score-MRI requires to perform nearly 150 sampling steps to process one scan and takes nearly 5 minutes with a single RTX5000 GPU, whereas our method takes only about 25 seconds per scan. The SMUG schemes also substantially outperform the deep equilibriumquilibrium model in the presence of perturbations.
In Fig 7 and Fig 8, we report PSNR and SSIM results of different methods at two sampling acceleration factors for the knee dataset. Therein, we observe quite similar outcomes to those reported in Fig 5.
Figs. 9 and 10 show reconstructed images by different methods for knee scans at 4x and 8x undersampling, respectively. We observe that SMUG and Weighted SMUG show fewer artifacts, sharper features, and fewer errors when compared to Vanilla MoDL and other baselines in the presence of the worst-case perturbations.
Fig. 11 presents average PSNR results over the test dataset for the considered models under different levels of worst-case perturbations (i.e., attack strength $\epsilon$). We used the knee dataset for this experiment. We observe that SMUG and weighted SMUG outperform RS-E2E, vanilla MoDL, and Deep-Eq across all perturbation strengths. When compared to Score-MRI and AT, our proposed methods consistently maintain higher PSNR values for moderate to large perturbations (less than $\epsilon=0.08$). For instance, when $\epsilon=0.02$, weighted SMUG reports more than 1 dB improvement over AT and Score-MRI.
Figure 11: PSNR of baseline methods and the proposed method versus perturbation strength (i.e., scaling) $\epsilon$ used in PGD-generated worst-case examples at testing time with 4x k-space undersampling. $\epsilon=0$ corresponds to clean accuracy.
IV-C Robustness for Varying Sampling Rates and Unrolling Steps
In this subsection, we evaluate the robustness of our proposed approaches and the considered baselines at varying sampling rates and unrolling steps.
For our first experiment, during training, a k-space undersampling or acceleration factor of 4x is used for our methods and the considered baselines. At testing time, we evaluate performance (in terms of PSNR) with acceleration factors ranging from 2x to 8x. The results are presented in Fig. 12. It is clear that when the acceleration factor during testing matches that of the training phase (4x), all methods achieve their highest PSNR results. Conversely, performance generally declines when the acceleration factors differ. For acceleration factors 3x to 8x (ignoring 4x where models were trained), we observe that our methods outperform all the considered baselines. For the 2x case, our methods report higher PSNR values compared to RS-E2E, vanilla MoDL, and Deep-Eq and slightly underperform AT, while Score-MRI shows more resilience at 2x.
For the second experiment, we study the performance of varying unrolling steps. More specifically, during training, we utilize 8 unrolling steps to train our methods and the baselines. At testing time, we report the results of utilizing 1 to 16 unrolling steps. The PSNR results of all the considered cases are given in Fig.  13. The results show that both SMUG and Weighted SMUG maintain performance comparable to the deep equilibriumquilibrium model. Furthermore, when using different unrolling steps and faced with additive measurement perturbations, the SMUG methods’ PSNR values are stable and close to the unperturbed case (indicating robustness), whereas the other methods see more drastic drop in performance. This behavior for SMUG also agrees with the theoretical bounds in Section III.
Although we do not intentionally design our method to mitigate MoDL’s instabilities against different sampling rates and unrolling steps, the SMUG approaches nevertheless provide improved PSNRs over other baselines. This indicates broader value for the robustification strategies incorporated in our schemes.
Figure 12: PSNR results for different MRI reconstruction methods versus different measurement sampling rates (models trained at 4×4$\times$4 × acceleration).
Figure 13: PSNR results for different MRI reconstruction methods at 4x k-space undersampling versus number of unrolling steps ($8$ steps used in training). “Clean” and ”Robust” denote the cases without and with added worst-case (for each method) measurement perturbations.
IV-D Importance of the Ustab Loss
We conduct additional studies on the unrolled stability loss in our scheme to show the importance of integrating target image denoising into SMUG’s training pipeline in (7). Fig. 14 presents PSNR values versus perturbation strength/scaling ($\epsilon$) when using different alternatives to $\mathcal{D}_{\bm{\theta}}(\mathbf{t})$ in (7), including $\mathbf{t}$ (the original target image), $\mathcal{D}_{\bm{\theta}}(\mathbf{x}_{n})$ (denoised output of each unrolling step), and variants when using the fixed, vanilla MoDL’s denoiser $\mathcal{D}_{\bm{\theta}_{\text{{MoDL}}}}$ instead. As we can see, the performance of SMUG varies when the UStab loss (7) is configured differently. The proposed $\mathcal{D}_{\bm{\theta}}(\mathbf{t})$ outperforms other baselines. A possible reason is that it infuses supervision of target images in an adaptive, denoising-friendly manner, i.e., taking the influence of $\mathcal{D}_{\bm{\theta}}$ into consideration.
Figure 14: PSNR vs. worst-case perturbation strength ($\epsilon)$ for SMUG for different configurations of UStab loss (7).
Refer to caption Refer to caption
Figure 15: Left: Norm of difference between SMUG and RS-E2E reconstructions and the ground truth for different choices of $\sigma$ in the smoothing process. A worst-case PGD perturbation $\bm{\delta}$ computed at $\epsilon=0.01$ was added to the measurements in all cases. Right: Robustness error for SMUG and RS-E2E at various $\sigma$, i.e., norm of difference between output with the perturbation $\bm{\delta}$ and without it.
Figure 16: Weights predicted by the weight encoder network in Weighted SMUG (from final layer of unrolling) plotted against root mean squared error (RMSE) of the corresponding denoised images for $5$ randomly selected scans (with 4x undersampling).
Figure 17: Reconstruction accuracy box plots for the fastMRI knee dataset with 4x acceleration factor for the case of ISTA-Net. The additive random Gaussian noise in the second column plots is obtained using a standard deviation of $0.01$. The worst-case additive noise in the third column is obtained using the PGD method with $\epsilon=0.02$.
IV-E Impact of the Noise Smoothing
To comprehensively assess the influence of the introduced noise during smoothing, denoted as $\bm{\eta}$, on the efficacy of the suggested approaches, we undertake an experiment involving varying noise standard deviations $\sigma$. The outcomes, documented in terms of RMSE, are showcased in Fig.15. The accuracy (reconstruction quality w.r.t. ground truth) and robustness error (error between with and without measurement perturbation cases) are shown for both SMUG and RS-E2E. We notice a notable trend: as the noise level $\sigma$ increases, the accuracy for both methods improves before beginning to degrade. Importantly, SMUG consistently outperforms end-to-end smoothing. Furthermore, the robustness error continually drops as $\sigma$ increases (corroborating with our analysis/bound in Section III), with more rapid decrease for SMUG.
IV-F Empirical Analysis of the behavior of Weighted SMUG
In subsequent final study, we analyze the behavior of the Weighted SMUG algorithm. We delve into the nuances of weighted smoothing, which can assign different weights to different images during the smoothing process. The aim is to gauge how the superior performance of Weighted SMUG arises from the variations in learned weights. Our findings indicate that among the $10$ Monte Carlo samplings implemented for the smoothing operation, those with lower denoising RMSE when compared to the ground truth images generally receive higher weights, as illustrated in Fig. 16.
IV-G Results of Applying Our Methods to ISTA-NET
In our concluding study, we investigate whether our robustification methods can be effective with an alternative unrolling technique, ISTA-Net [14]. For ISTA-Net, we adopted the default architecture, utilizing the ADAM optimizer with a learning rate of $10^{-4}$. The network was configured with $9$ phases (unrolling iterations) and trained on the fastMRI knee dataset comprising $3000$ scans at 4x undersampling and with $100$ epochs for training. Similar to previous experiments, we used $64$ scans for testing. Other settings for training the vanilla ISTA-Net were set to default values [^2] https://github.com/jianzhangcs/ISTA-Net-PyTorch. Other settings for the RS-E2E version, and the SMUG and Weighted SMUG versions of ISTA-Net were similar to the MoDL case. The results, as presented in Figure 17, demonstrate that the clean accuracy performance of SMUG and weighted SMUG versions of ISTA-Net are comparable to vanilla ISTA-Net. Notably, under conditions of random noise (Gaussian noise with $\sigma=0.01$ added) and PGD attack (30 steps with $\epsilon=0.02$) perturbed measurements, our method surpasses both the original ISTA-Net and the RS-E2E version. The comparative results reveal that the performance closely aligns with the outcomes previously observed when unrolling smoothing was combined with the MoDL network.
V Discussion and Conclusion
In this work, we proposed a scheme for improving the robustness of DL-based MRI reconstruction. In particular, we investigated deep unrolled reconstruction’s weaknesses in robustness against worst-case or noise-like additive perturbations, sampling rates, and unrolling steps. To improve the robustness of the unrolled scheme, we proposed SMUG with a novel unrolled smoothing loss. We also provided a theoretical analysis on the robustness achieved by our proposed method. Compared to the vanilla MoDL approach and other schemes, we empirically showed that our approach is effective and can significantly improve the robustness of a deep unrolled scheme against a diverse set of external perturbations. We also further improved SMUG’s robustness by introducing weighted smoothing as an alternative to conventional RS, which adaptively weights different images when aggregating them. In future work, we hope to apply the proposed schemes to other imaging modalities and evaluate robustness against additional types of realistic perturbations. While we theoretically characterized the robustness error for SMUG, we hope to further analyze its accuracy-robustness trade-off with perturbations.
Appendix A Proof of Theorem 1
A-A Preliminary of Theorem 1
Lemma 1.
Let $f:\mathbb{R}^{d}\to\mathbb{R}^{m}$ be any bounded function. Let $\bm{\eta}\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})$. We define $g:\mathbb{R}^{d}\to\mathbb{R}^{m}$ as
$g(\mathbf{x})=\mathbb{E}_{\bm{\eta}}[f(\mathbf{x}+\bm{\eta})].$
Then, $g$ is an $\frac{M}{\sqrt{2\pi}\sigma}$-Lipschitz map, where $M=2\max_{\mathbf{x}\in\mathbb{R}^{d}}(\|f(\mathbf{x})\|_{2})$. In particular, for any $\mathbf{x},\bm{\delta}\in\mathbb{R}^{d}$:
$\|g(\mathbf{x})-g(\mathbf{x}+\bm{\delta})\|_{2}\leq\frac{M}{\sqrt{2\pi}\sigma}\|\bm{\delta}\|_{2}.$
Proof.
[…]
which is a standard result for the norm of an integral. We further apply Holder’s inequality to upper bound $\|g(\mathbf{x})-g(\mathbf{x}+\bm{\delta})\|_{2}$ with
[…]
Observe that $\mu(\bm{w}-\mathbf{x})\geq\mu(\bm{w}-\mathbf{x}-\bm{\delta})$ if $\|\bm{w}-\mathbf{x}\|_{2}\leq\|\bm{w}-\mathbf{x}-\bm{\delta}\|_{2}$. Let $D=\{\bm{w}:\|\bm{w}-\mathbf{x}\|_{2}\leq\|\bm{w}-\mathbf{x}-\bm{\delta}\|_{2}\}$. Then, we can rewrite the above bound as
[…]
Following Lemma 3 in [46], we obtain the bound
[…]
which implies that $\|g(\mathbf{x})-g(\mathbf{x}+\bm{\delta})\|_{2}\leq\frac{2\max_{\mathbf{x}\in\mathbb{R}^{d}}(\|f(\mathbf{x})\|_{2})}{\sqrt{2\pi}\sigma}\|\bm{\delta}\|_{2}=\frac{M}{\sqrt{2\pi}\sigma}\|\bm{\delta}\|_{2}$. This completes the proof. ∎
A-B Proof of Theorem 1
Proof.
Assume that the data consistency step in MoDL at iteration $n$ is denoted by $\mathbf{x}^{n}_{\text{M}}(\mathbf{A}^{H}\mathbf{y})$. We will sometimes drop the input and $\mathbf{y}$ dependence for notational simplicity. Then
$\mathbf{x}^{1}_{\text{M}}=(\mathbf{A}^{H}\mathbf{A}+\mathbf{I})^{-1}(\mathbf{A}^{H}\mathbf{y}+\mathcal{D}_{\bm{\theta}}(\mathbf{A}^{H}\mathbf{y}))\>,$ (21)
$\mathbf{x}^{n}_{\text{M}}=(\mathbf{A}^{H}\mathbf{A}+\mathbf{I})^{-1}(\mathbf{A}^{H}\mathbf{y}+\mathcal{D}_{\bm{\theta}}(\mathbf{x}^{n-1}_{\text{M}})),$ (22)
where $\mathcal{D}_{\bm{\theta}}$ is the denoiser function. For the sake of simplicity and consistency with the experiments, we use the weighting parameter $\lambda=1$ (in the data consistency step). We note that the proof works for arbitrary $\lambda$. SMUG introduces an iteration-wise smoothing step into MoDL as follows:
$\mathbf{x}_{\text{S}}^{1}=((\mathbf{A}^{H}\mathbf{A}+\mathbf{I})^{-1}(\mathbf{A}^{H}\mathbf{y}+\mathbb{E}_{\bm{\eta}_{1}}[\mathcal{D}_{\bm{\theta}}(\mathbf{A}^{H}\mathbf{y}+\bm{\eta}_{1})])$ (23)
$\mathbf{x}_{\text{S}}^{n}=((\mathbf{A}^{H}\mathbf{A}+\mathbf{I})^{-1}(\mathbf{A}^{H}\mathbf{y}+\mathbb{E}_{\bm{\eta}_{n}}[\mathcal{D}_{\bm{\theta}}(\mathbf{x}_{\text{S}}^{n-1}+\bm{\eta}_{n})])$ (24)
$=(\mathbf{A}^{H}\mathbf{A}+\mathbf{I})^{-1}(\mathbf{A}^{H}\mathbf{y})+$ (25)
$(\mathbf{A}^{H}\mathbf{A}+\mathbf{I})^{-1}\mathbb{E}_{\bm{\eta}_{n}}[\mathcal{D}_{\bm{\theta}}(\mathbf{x}_{\text{S}}^{n-1}+\bm{\eta}_{n})],$
where we apply the expectation to the denoiser $\mathcal{D}_{\bm{\theta}}$ at each iteration. We use $\bm{\eta}_{n}$ to denote the noise during smoothing at iteration $n$. The robustness error of SMUG after $n$ iterations is $\|\mathbf{x}_{\text{S}}^{n}(\mathbf{A}^{H}\mathbf{y})-\mathbf{x}_{\text{S}}^{n}(\mathbf{A}^{H}(\mathbf{y}+\bm{\delta}))\|$. We apply Lemma 1 and properties of the norm (e.g., triangle inequality) to bound $\|\mathbf{x}_{\text{S}}^{n}(\mathbf{A}^{H}\mathbf{y})-\mathbf{x}_{\text{S}}^{n}(\mathbf{A}^{H}(\mathbf{y}+\bm{\mathbf{\delta}}))\|$ as
$\leq\|(\mathbf{A}^{H}\mathbf{A}+\mathbf{I})^{-1}\mathbf{A}^{H}\bm{\delta}\|$ (26)
$+\|(\mathbf{A}^{H}\mathbf{A}+\mathbf{I})^{-1}\cdot\bigl{(}\mathbb{E}_{\bm{\eta}_{n}}[\mathcal{D}_{\bm{\theta}}\bigl{(}\mathbf{x}_{\text{S}}^{n-1}(\mathbf{A}^{H}\mathbf{y})+\bm{\eta}_{n}\bigr{)}]-$
$\mathbb{E}_{\bm{\eta}_{n}}[\mathcal{D}_{\bm{\theta}}\bigl{(}\mathbf{x}_{\text{S}}^{n-1}(\mathbf{A}^{H}(\mathbf{y}+\bm{\delta}))+\bm{\eta}_{n}\bigr{)}]\bigr{)}\|$
$\leq\|(\mathbf{A}^{H}\mathbf{A}+\mathbf{I})^{-1}\|_{2}\|\mathbf{A}^{H}\bm{\delta}\|_{2}$
$+\|(\mathbf{A}^{H}\mathbf{A}+\mathbf{I})^{-1}\|_{2}\|\mathbb{E}_{\bm{\eta}_{n}}[\mathcal{D}_{\bm{\theta}}\bigl{(}\mathbf{x}_{\text{S}}^{n-1}(\mathbf{A}^{H}\mathbf{y})+\bm{\eta}_{n}\bigr{)}]-$
$\mathbb{E}_{\bm{\eta}_{n}}[\mathcal{D}_{\bm{\theta}}\bigl{(}\mathbf{x}_{\text{S}}^{n-1}(\mathbf{A}^{H}(\mathbf{y}+\bm{\delta}))+\bm{\eta}_{n}\bigr{)}]\|$
$\leq\|(\mathbf{A}^{H}\mathbf{A}+\mathbf{I})^{-1}\|_{2}\|\mathbf{A}^{H}\bm{\delta}\|_{2}+\|(\mathbf{A}^{H}\mathbf{A}+\mathbf{I})^{-1}\|_{2}\times$ (27)
$\begin{pmatrix}\frac{M}{\sqrt{2\pi}\sigma}\end{pmatrix}\|\mathbf{x}_{\text{S}}^{n-1}(\mathbf{A}^{H}\mathbf{y})-\mathbf{x}_{\text{S}}^{n-1}(\mathbf{A}^{H}(\mathbf{y}+\bm{\mathbf{\delta}}))\|.$
Here, $M=2\max_{\mathbf{x}}(\|\mathcal{D}_{\bm{\theta}}(\mathbf{x})\|)$. Then we plug in the expressions for $\mathbf{x}_{\text{S}}^{n-1}(\mathbf{A}^{H}\mathbf{y})$ and $\mathbf{x}_{\text{S}}^{n-1}(\mathbf{A}^{H}(\mathbf{y}+\bm{\mathbf{\delta}}))$ (from (24)) and bound their normed difference with $\|(\mathbf{A}^{H}\mathbf{A}+\mathbf{I})^{-1}\mathbf{A}^{H}\bm{\delta}\|+\|(\mathbf{A}^{H}\mathbf{A}+\mathbf{I})^{-1}\cdot\bigl{(}\mathbb{E}_{\bm{\eta}_{n-1}}[\mathcal{D}_{\bm{\theta}}\bigl{(}\mathbf{x}_{\text{S}}^{n-2}(\mathbf{A}^{H}\mathbf{y})+\bm{\eta}_{n-1}\bigr{)}]-\mathbb{E}_{\bm{\eta}_{n-1}}[\mathcal{D}_{\bm{\theta}}\bigl{(}\mathbf{x}_{\text{S}}^{n-2}(\mathbf{A}^{H}(\mathbf{y}+\bm{\delta}))+\bm{\eta}_{n-1}\bigr{)}]\bigr{)}\|$. This is bounded above similarly as for (26). We repeat this process until we reach the initial $\mathbf{x}_{\text{S}}^{0}$ on the right hand side. This yields the following bound involving a geometric series.
$\|\mathbf{x}_{\text{S}}^{n}(\mathbf{A}^{H}\mathbf{y})-\mathbf{x}_{\text{S}}^{n}(\mathbf{A}^{H}(\mathbf{y}+\bm{\mathbf{\delta}}))\|$ (28)
$\leq\|\mathbf{A}^{H}\bm{\delta}\|_{2}\begin{pmatrix}\sum_{j=1}^{n}\|(\mathbf{A}^{H}\mathbf{A}+\mathbf{I})^{-1}\|_{2}^{j}\cdot\begin{pmatrix}\frac{M}{\sqrt{2\pi}\sigma}\end{pmatrix}^{j-1}\end{pmatrix}$
$+\|(\mathbf{A}^{H}\mathbf{A}+\mathbf{I})^{-1}\|_{2}^{n}\begin{pmatrix}\frac{M}{\sqrt{2\pi}\sigma}\end{pmatrix}^{n}\|\mathbf{A}^{H}\bm{\delta}\|_{2}$ (29)
$\leq\|\mathbf{A}\|_{2}\|\bm{\delta}\|_{2}\|(\mathbf{A}^{H}\mathbf{A}+\mathbf{I})^{-1}\|_{2}\begin{pmatrix}\frac{1-\begin{pmatrix}\frac{M}{\sqrt{2\pi}\sigma}\end{pmatrix}^{n}\|(\mathbf{A}^{H}\mathbf{A}+\mathbf{I})^{-1}\|_{2}^{n}}{1-\frac{M}{\sqrt{2\pi}\sigma}\|(\mathbf{A}^{H}\mathbf{A}+\mathbf{I})^{-1}\|_{2}}\end{pmatrix}$
$+\|(\mathbf{A}^{H}\mathbf{A}+\mathbf{I})^{-1}\|_{2}^{n}\begin{pmatrix}\frac{M}{\sqrt{2\pi}\sigma}\end{pmatrix}^{n}\|\mathbf{A}\|_{2}\|\bm{\delta}\|_{2}\leq C_{n}\|\bm{\delta}\|_{2},$ (30)
where we used the geometric series formula, and $C_{n}=\alpha\|\mathbf{A}\|_{2}\begin{pmatrix}\frac{1-\begin{pmatrix}\frac{M\alpha}{\sqrt{2\pi}\sigma}\end{pmatrix}^{n}}{1-\frac{M\alpha}{\sqrt{2\pi}\sigma}}\end{pmatrix}+\|\mathbf{A}\|_{2}\begin{pmatrix}\frac{M\alpha}{\sqrt{2\pi}\sigma}\end{pmatrix}^{n}$, with $\alpha=\|(\mathbf{A}^{H}\mathbf{A}+\mathbf{I})^{-1}\|_{2}$. ∎